OpenStudy (hulkart):

In a certain lottery, five different numbers between 1 and 20 inclusive are drawn. To win the lottery, a person must select the correct 5 numbers in the same order in which they were drawn. What is the probability of winning?

OpenStudy (kropot72):

To solve this question you need to find the number of permutations of the 20 numbers taken 5 at a time: \[\large P(20, 5)=\frac{20!}{(20-5)!}=you\ can\ calculate\] When you have this number, the inverse of it is the solution.

OpenStudy (hulkart):

just give the answer please??

OpenStudy (academicgurusinc):

We are not supposed to give out answers. But rather guide you and help you get there. Well your chance of getting the first right is 1/20 then 1/19 and so on until 1/16 for the fifth ball. Here's why: So 1/(20*19*18*17*16) = 1/860,480 Another way to write the answer is 1/P(20,5) or 1/20P5 where nPk = P(n,k) = C(n,k) * k! and C(n,k) is binomial coefficient and k! is factorial. Hope that helps. If you found this helpful, I encourage to subscribe to our youtube channel https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)